# Why Static Friction Decreases From Single to Multi-asperity Contacts

### v1

@@ Abstract
In conclusion, we have investigated how the macroscopic frictional response of a Si-on-Si system evolves from a single to a multi-asperity contact interface, and shown how the macroscopic static friction coefficient decreases with increasing normal load across a broad range of loads. We propose a stick-slip model, analogous to a fiber bundle model, based on elastic contact calculations to capture the decrease in macroscopic static friction coefficient and friction drop with increasing normal load. This decrease is found to be dominated by the relative fraction of critical contact asperities at the interface that collectively fail when the maximal static friction force is reached. With increasing normal load, we observe the growing populations of pre-sliding asperities -contact asperities that slip before the maximal tangential force is applied to the interface- and subcritical asperities -contact asperities whose slip is activated by redistribution of the excess friction from critical asperities. Both the pre-sliding and subcritical populations of asperities observed at the critical interface state grow with increasing normal load, causing the static friction coefficient to drop with increased normal force. By changing the hold time in our hold-slid-hold experiments, we demonstrated that these results remain valid when considering a different reference nano-scale friction behavior. While our experiments were limited to Si-on-Si interface, we believe our results to be relevant for the transition from single to multi-asperity contact of generic interfaces, for which the elastic interactions between contact asperities remain negligible. We expect the range of normal load over which the decrease in macroscopic friction occurs to be dependent on both the material properties and the details of the surface (e.g., its roughness), an issue that should be addressed in future work. All in all, our findings shed new light on the complexity that emerges when contact interfaces scale up from single asperity to multi-asperity - a central question in tribology and an important challenge in precision positioning.
@@ Acknowledgements.
We thank R. Carpick, B. Mcclimon, J. Dijksman and K. Farain for valuable discussions. TR acknowledges support of the Swiss National Science Foundation through the fellowship No. P500PT/222342. LP thanks the invaluable support from Min Li.
@@
Supplemental Material
Figure S1: AFM topography. (a)(b)(c) Three different AFM topographies measured from silicon balls used in experiments. Measurements were taken using tapping mode in AFM (Dimension Icon, Bruker) with Si tips (RTESPA-300, Bruker). The topography size is 31.13 × 31.13 $\mu\text{m}^{2}$ with a RMS (root-mean-square) roughness around 40 nm. Scale bar, 5 $\mu$m.
Figure S2: Nanoscale friction coefficient-displacement curve used for simulation in Fig. 2. The black and red curves correspond to the AFM SHS measurements for a hold time of 100 s and 10 s, respectively. The static and dynamic friction values for a hold time of 100 s are indicated respectively with $\mu^{S}$ and $\mu^{D}$.
@@
Figure S3: Topography of silicon ball apex after friction experiments. (a) Optical image of the apex of silicon ball captured over an area of 283 $\times$ 212 $\mu$m2, using a laser-scanning confocal profilometer (Keyence VK-X1000). (b) Height profile derived from the optical image in (a). (c) AFM topography with a size of 90 $\times$ 90 $\mu$m2, which corresponds to the square-marked area in (a) and (c). Both optical and AFM images reveal no visible signs of wear, indicating minimal impacts of wear on our experimental result. Scale bar, 50 $\mu$m.
Supplemental movies
@@
Each movie shows the evolution of the distribution of local friction coefficients with increasing imposed displacement, with the critical asperities ($\mu=\mu^{S}$) in brown, the pre-sliding asperities ($\mu=\mu^{D}$) in yellow, and the subcritical asperities ($\mu<\mu^{0}$) in green. The movies are available online.

### v2

@@ Abstract
In conclusion, we have investigated how the macroscopic frictional response of a Si-on-Si system evolves from a single to a multi-asperity contact interface, and show how the macroscopic static friction coefficient decreases with increasing normal load across a broad range of loads. We propose a stick-slip model, analogous to a fiber bundle model, based on elastic contact calculations to capture the decrease in macroscopic static friction coefficient and friction drop with increasing normal load. This decrease is found to be dominated by the relative fraction of critical contact elements at the interface that collectively fail when the maximal static friction force is reached. With increasing normal load, we observe the growing populations of pre-sliding elements -contact elements that slip before the maximal tangential force is applied to the interface- and subcritical elements -contact elements whose slip is activated by redistribution of the excess friction from critical elements. Both the pre-sliding and subcritical populations of elements observed at the critical interface state grow with increasing normal load, causing the friction drop to decrease with increasing normal force. By changing the hold time in our hold-slide-hold experiments, we demonstrated that these results remain valid when considering a different reference nanoscale friction behavior. All in all, our findings shed new light on the complexity that emerges when contact interfaces scale up from single asperity to multi-asperity - a central question in tribology and an important challenge in precision positioning.
@@ Acknowledgements.
We thank R. Carpick, B. Mcclimon, J. Dijksman and K. Farain for valuable discussions. TR acknowledges support of the Swiss National Science Foundation through the fellowship No. P500PT/222342. LP acknowledges funding from the China Scholarship Council and thanks the invaluable support from Min Li.
@@
Appendix A Supplemental Material for: The Decrease of Static Friction Coefficient with Interface Growth from Single to Multi-asperity Contact
A Experimental Method
The rheometer (DSR 502, Anton Paar) and an analytical balance (Quintix, Satorious) were used to conduct the friction measurements. A 3 mm-diameter silicon ball was clamped to the geometry of a rheometer and brought into contact with a homemade Si cantilever that was made by gluing a piece of 1×1 cm p-type silicon (100) wafer (University Wafer) to a brass sheet. The normal stiffness of the Si cantilever could be tuned by changing the width and thickness of the used brass sheet. The Si cantilever was mounted on a stage, which was attached to the weighing pan of an analytical balance by dead weight. Before the friction experiments, the (naturally oxidized) silicon balls (Goodfellow) were sonicated in ethanol and then Milli-Q water, followed by nitrogen flow drying. To compare with the highly hydroxylated silicon surfaces used in previous AFM experiments [16, 15], the silicon balls and the silicon cantilevers were treated by oxygen plasma cleaning inside a reactor (Zepto One, Diener Electronic) [23]. In the friction experiments, both the plasma-cleaned silicon ball and the plasma-cleaned silicon cantilever were initially equilibrated for one hour inside the dry chamber before contact was formed between them. The relative humidity inside the chamber was stabilized at 0.8% by continuously flowing dry nitrogen. During sliding, the rheometer and balance simultaneously measured the friction force ($F_{f}$) and normal load ($N$) at a sampling rate of 20 Hz, corresponding to one data point for every 15 nm of displacement. When the sliding was restarted, we reported the peak static friction ($F_{S}$ or $\mu_{S}$) and the average value of the dynamic friction ($F_{D}$ or $\mu_{D}$) in the steady state sliding regime, measured over a displacement of $\sim 1\,\mu\text{m}$ after the peak friction. We show in Fig. S1 the static and dynamic friction coefficients at a hold time of 100 s.
Figure S1: Friction coefficient as a function of normal load at a hold time of 100 s. The black and red data correspond to the static and dynamic friction coefficient, respectively. The measured static friction coefficient shows a decreasing trend with increasing normal load, while the dynamic friction coefficient remains relatively constant under the same conditions. Notably, a dip in both static and dynamic friction coefficients is observed around a normal load of 1 mN, which does not affect the friction drop reported in the main text, as both coefficients contribute equally to the ratio. We suspect that the dip is caused by a resonance in the system, as oscillating normal force is known to decrease friction [33, 49]. The contact stiffness decreases with decreasing normal force and may enable vibrations in the normal direction, thereby lowering both static and dynamic friction, at around 1 mN.
The application of shear force from above (and below) the interface can cause torque in friction experiments, which may impact the distribution of normal stress across the interface. In flat-on-flat geometries, this effect has been utilized to ensure that the onset of failure is located at the leading or trailing edge of the interface [35]. In the sphere-on-flat experimental geometry we use here, the redistribution of the interfacial normal stress due to torque is negligible [8].
@@ A Experimental Method
Figure S2: Topography of silicon ball apex after friction experiments. (a) Optical image of the apex of silicon ball captured over an area of 283 $\times$ 212 $\mu$m2, using a laser-scanning confocal profilometer (Keyence VK-X1000). (b) The height profile derived from the optical image in (a). (c) AFM topography with a size of 90 $\times$ 90 $\mu$m2, which corresponds to the square-marked area in (a) and (c). Both optical and AFM images reveal no visible signs of wear, indicating the minimal impact of wear on our experimental result. Scale bar, 50 $\mu$m.
Figure S3: AFM topography. (a)(b)(c) Three different AFM topographies measured from silicon balls used in experiments. Measurements were taken using tapping mode in AFM (Dimension Icon, Bruker) with Si tips (RTESPA-300, Bruker). The topography size is 31.13 × 31.13 $\mu\text{m}^{2}$ with an RMS (root-mean-square) roughness around 40 nm. Scale bar, 5 $\mu$m..
B Numerical Method
We use an open-source implementation of the boundary integral method [38] for the elastic contact calculation. The contact between a silicon ball and a silicon wafer is modeled as a single rough surface in contact with a rigid flat surface, with the roughness from both silicon surfaces mapped onto the single rough surface. The contact problem was then solved with a compound roughness $h=h_{1}-h_{2}$, with $h_{1}$ the roughness of the silicon ball and $h_{2}$ the roughness of the silicon wafer measured with AFM after friction experiments, see Fig. S2 and Fig. S3. The surface topography size is 31.13 × 31.13 $\mu\text{m}^{2}$ and is made of 1500 × 1500 elements, each with a length of 20.75 × 20.75 nm. We use the mechanical properties of silicon: Young’s modulus: $E=130\pm 3\,\text{GPa}$, Poisson’s ratio: $\nu=0.26$. The apparent pressure is imposed and varied to get the contact pressure map at various total normal loads. Examples of the contact maps under various loads are presented in Fig. S4, illustrating the evolution from single to multi-contact asperity interface. For $N=0.1$mN, the contact interface is made of three continuous contact patches (each made of several elements), see the inset in Fig. S4(a). The total contact area in this case spans around 50 elements. Two of these contact patches eventually merge to form a single contact when the normal load is increased to $N=1$mN, accompanied by the formation of a new contact patch at another location of the interface, see Fig. S4(b). The map of the contact under a load of $N=10$mN (Fig. S4(c)) is radically different, with the contact made of numerous contact patches (each of them several elements in size).
Figure S4: Contact maps under normal loads of (a) $N=0.1$mN, (b) $N=1$mN, (c) $N=10$mN and (d) $N=100$mN. The inset in panel (a) zooms into the only area exhibiting contact under this load. The insets in panel (b) show the two main areas of contact.
For the stick-slip model, the friction behavior at the element level follows the measurements from AFM experiments [15]. The behavior for 100 s and 10 s waiting times are shown in Fig. S5.
Figure S5: Nanoscale friction coefficient-displacement curve used for simulation in Fig. 2. The black and red curves correspond to the AFM SHS measurements for a hold time of 100 s and 10 s, respectively. The static and dynamic friction values for a hold time of 100 s are indicated respectively with $\mu^{S}$ and $\mu^{D}$. The nanoscale friction coefficient is calculated as the ratio of the friction force to the total normal force, comprising the adhesion force and the externally applied load.
C Supplemental movies
@@
Each movie shows the evolution of the distribution of local stress ratios with increasing imposed displacement, with the critical asperities ($\mu=\mu^{S}$) in brown, the pre-sliding asperities ($\mu=\mu^{D}$) in yellow, and the subcritical asperities ($\mu<\mu^{0}$) in green. The movies are available online. Snapshots of the movies are shown in Fig.S6.
[…]